3. Assume an economy in which only broccoli and cauliflower are produced. In year 1, 300 million pounds of broccoli are produced and consumed, and its price is S0.50 per pound, while 400 million pounds of cauliflower are produced and consumed and its price is S0.40 per pound. In year 2 250 million pounds of broccoli are produced and consumed and its price is S0.50 per pound, while 400 million pounds of cauliflower are produced and its price is SO.80 per pound (a) Using year 1 as the base year, calculate the GDP price deflator in years 1 and 2, and calculate the rate of inflation between years 1 and 2 from the GDP price deflator (b) Using year 1 as the base year, calculate the CPI in years 1 and 2, and calculate the CPI rate of inflation. Explain any differences in your results between parts (a) and (b)
Answer #1

SOLUTION:

1)

Good

Quantity

Price

Quantity

Price

Broccoli

300

0.5

250

0.5

Cauliflower

400

0.4

400

0.8

Year 1 nominal GDP = 300* 0.5 + 400 * 0.4 = 310

Year 1 real GDP = 300* 0.5 + 400 * 0.4 = 310

Year 2 nominal GDP = 250 * 0.5 + 400 * 0.8 = 445

Year 2 real GDP = 250*0.5 + 400 * 0.4 = 285

Year 1 GDP deflator equals

Year 2 GDP deflator: 445/285 * 100 = 156.14%

The percentage change in the deflator is 56.14%

 

2) CPI Year 1 = 100

CPI Year 2 = (300 * 0.5 + 400 * 0.8) / (300 * 0.5 + 400 * 0.4) * 100 = 470/310 * 100 = 151.61

The percentage change in CPI is 51.61%